OpenStudy (anonymous):

Decide whether the formula an = –4n(n – 1) is explicit or recursive. Then find the first five terms of the sequence.

OpenStudy (amistre64):

do you see if there is a common difference between terms?

OpenStudy (amistre64):

does -7-(-5.6) = -4.2 - (-2.8) by chance?

OpenStudy (amistre64):

if there is a common difference, then we can write up an arithmetic progression of the form: an = a1+d(n-1)

OpenStudy (amistre64):

or recuursively as: an = a(n-1) + d